# From Source to Target: Leveraging Transfer Learning for Predictive Process Monitoring in Organizations

### v1

@@
Event logs reflect the behavior of business processes that are mapped in organizational information systems. Predictive process monitoring (PPM) transforms these data into value by creating process-related predictions that provide the insights required for proactive interventions at process runtime. Existing PPM techniques require sufficient amounts of event data or other relevant resources that might not be readily available, preventing some organizations from utilizing PPM. The transfer learning-based PPM technique presented in this paper allows organizations without suitable event data or other relevant resources to implement PPM for effective decision support. The technique is instantiated in two real-life use cases, based on which numerical experiments are performed using event logs for IT service management processes in an intra- and inter-organizational setting. The results of the experiments suggest that knowledge of one business process can be transferred to a similar business process in the same or a different organization to enable effective PPM in the target context. With the proposed technique, organizations can benefit from transfer learning in an intra- and inter-organizational setting, where resources like pre-trained models are transferred within and across organizational boundaries.
@@ 1 Introduction
Business processes are the backbone of organizational value creation (Dumas et al. 2018). With ongoing digitalization, business processes are increasingly mapped in information systems (IS), such as enterprise resource planning (ERP) systems, and executed by process users and other stakeholders (Beverungen et al. 2021). Thus, business processes leave large amounts of digital event data reflecting their behavior (van der Aalst 2016). To transform these data into value, various business process analytics approaches have been developed (zur Muehlen and Shapiro 2009).
One of these approaches is predictive process monitoring (PPM) – a set of techniques that leverage historical event log data (Maggi et al. 2014) to create prediction models Grigori et al. (2004), capable of predicting various targets in running business process instances, such as the next activity, remaining time, or process-related outcomes (Di Francescomarino et al. 2018). By generating these process-related predictions, PPM enables proactive interventions in running business process instances, helping to reduce risks and optimize resource allocation (Di Francescomarino et al. 2018). This leads to improved business process performance (Marquez-Chamorro et al. 2017), even across organizational boundaries (Oberdorf et al. 2023). Therefore, PPM is gaining momentum in business process management (BPM), while its techniques are considered a new class of process mining techniques (Di Francescomarino and Ghidini 2022).
Most PPM techniques rely on shallow machine learning (ML) models (Marquez-Chamorro et al. 2017) or deep learning (DL) models (Rama-Maneiro et al. 2023) to generalize knowledge from historical to unseen event log data (Weinzierl et al. 2024). These models are trained on historical event log data and then applied to new instances of a running business process in order to predict various targets depending on the business context and analytical goals (Maggi et al. 2014). However, organizations – or different departments or subsidiaries within the same organization – often lack the prerequisites for building effective prediction models. These include access to sufficient amounts of event log data (Käppel et al. 2021) and the technical infrastructure required to train complex models, such as graphical processing units.
@@
The field of transfer learning (TL) provides techniques specifically designed to support such scenarios (Pan and Yang 2009), enabling the transfer of ML models and the knowledge they contain from one learning problem (i.e., domain and task) to another. In PPM, TL can be applied in inter-organizational scenarios, in which prediction models are transferred between processes of different organizations, and in intra-organizational scenarios, where a prediction model is transferred from one process to another within a single organization (e.g., across departments or subsidiaries). In both scenarios, we consider all resources to be transferred as process knowledge. This includes the prediction model (also called the base model). Process knowledge originates in its source context, which includes the business process itself, the event data generated by it, and the specific PPM task being addressed using the event data of the business process. The transferred process knowledge is then applied in the target context (see Figure 1).
Figure 1: Transfer process knowledge from source to target contexts in PPM.
Building on this idea, we present a TL-based technique for PPM that relies on existing similarities between business processes and aims to apply a prediction model to a target context. Therefore, the prediction model is transferred together with further resources, such as information on the corresponding encoding strategy. In particular, this work contributes to the existing body of knowledge in four ways:
@@
    We present a TL-based technique for PPM that enables the transfer of a base model for an outcome-oriented prediction target without the necessity to fine-tune it on the event data of the target context. Our approach also provides guidance on which resources need to be transferred, including pre-trained models, information on the corresponding encoding strategy, further preprocessing details, and metrics used for measuring the prediction performance.
@@
    We propose a cross-domain encoding strategy for event data that enables effective knowledge transfer from source to target contexts. While timestamp information is encoded via a relative mapping approach, activity information is encoded using pre-trained embedding models to preserve semantic consistency across contexts.
@@
    We empirically evaluate our TL-based PPM technique in both an intra- and inter-organizational ITSM use case. The results show that the transferred base model achieves a clearly higher prediction performance than models built with traditional PPM techniques on the event data of the target context.
To carry out our research, we apply computational design science research (DSR) (Rai et al. 2017) and structure our work according to the DSR publication schema by Gregor and Hevner (2013): First, we present relevant background on ML and DL in PPM, and TL in general (Section 2). Second, we explain the practical relevance of TL for PPM in organizations and introduce inter- and intra-organizational TL in the context of PPM (Section 3). Subsequently, we give an overview of related work on TL in PPM and position our proposed technique with regard to them (Section 4). Then, we provide a detailed description of the proposed artifact (Section 5). Afterward, the evaluation of our artifact is described by introducing the intra- and inter-organizational use cases, and providing details on the setup used in the evaluation (Section 6). Then, we present the results of our evaluation (Section 7), discuss implications and future research directions (Section 8), and conclude the paper (Section 9).
@@ 2.1 Event log data and predictive process monitoring
A business process is defined as a sequence of interrelated activities and decisions undertaken to achieve a valuable outcome for a customer (Dumas et al. 2018). Each execution of a business process is referred to as a process instance or case (van der Aalst 2016). Information technology (IT) systems (e.g., ERP systems) commonly record these executions in the form of event logs, that is, collections of timestamped events described by various event attributes that store information about the execution of activities (i.e., steps in the process) (van der Aalst 2016, p. 130). Each event contains at least a case identifier, the name of the executed activity, and a timestamp indicating when the event occurred. Events belonging to the same process instance can be temporally ordered by their timestamp to form a so-called trace (van der Aalst 2016, p. 134).
@@ 2.2 Machine learning in predictive process monitoring
A branch of early PPM approaches creates prediction models by augmenting discovered process models (or structurally similar representations such as Petri nets or finite automatons) with predictive capabilities (e.g., van der Aalst et al. 2011; Rogge-Solti and Weske 2015). However, as real-world processes are usually more complex than reflected in discovered process models, which frequently fail to capture all execution scenarios, their predictive power is limited (van der Aalst 2011). To overcome this limitation, a more recent branch of PPM approaches emerged that exploit the generalization capabilities of ML in identifying and generalizing patterns in event log data to learn prediction models in a data-driven manner.
@@
Besides a wide range of conventional ML algorithms such as linear regressions, logistic regressions, decision trees, random forests, and gradient boosted decision trees, deep neural networks (DNNs), from the field of DL, have been applied for building prediction models (Marquez-Chamorro et al. 2017; Teinemaa et al. 2019; Verenich et al. 2019). Especially, the latter has gained prominence due to its ability to learn complex patterns and dependencies in large datasets (LeCun et al. 2015). DL techniques have shown superior prediction performance and generalizability across various PPM tasks (Kratsch et al. 2021), and have thus become the state of the art.
@@
In recent years, a variety of neural network architectures have been employed for different prediction targets, including multi-layer perceptrons (MLPs), convolutional neural networks (CNNs), and recurrent neural network (RNN) with either LSTM cells or gated recurrent units (GRUs), with LSTM networks being favored by the majority (Neu et al. 2022). Among these, RNN, particularly LSTM networks, have been the most widely adopted in PPM. For example, Evermann et al. (2017) applied LSTMs for next activity prediction, while Tax et al. (2017) demonstrated that LSTMs outperform conventional ML approaches in predicting both the next activity and its corresponding timestamp. More recently, transformer-based models, originally introduced in natural language processing, have shown high prediction performance. These models utilize self-attention mechanisms to effectively capture long-range dependencies in sequential data, as present in event data. For instance, Zaharah A. Bukhsh (2021) proposed a transformer model capable of predicting next activities, remaining time, and timestamps. Beyond efforts to enhance prediction performance through increasingly sophisticated and powerful neural network architectures, complementary strategies have focused on enriching the input data by incorporating additional sources of information. While many prediction models primarily rely on control-flow information and time information derived from timestamps (e.g., elapsed time), several approaches extend this by integrating contextual information, for example, in the form of case-level or resource attributes (Weinzierl et al. 2019). For example, Di Francescomarino et al. (2018) proposed multi-perspective models that combine control flow and contextual information for outcome prediction. As another example, Hinkka et al. (2019) suggested clustering categorical event attributes and using the resulting cluster labels as additional information in a RNN. In contrast to these approaches, Pegoraro et al. (2021) and Cabrera et al. (2022) integrate unstructured data in the form of text as contextual information. The ability of DL methods to handle diverse prediction tasks and to incorporate heterogeneous data sources not only demonstrates their flexibility but also hints at the potential for reusing learned representations in varying contexts (Weinzierl et al. 2024).
@@
To build a prediction model using an ML or DL algorithm, it is crucial to appropriately encode the underlying event data, as the quality and structure of the encoding strongly affect the model’s ability to generalize. Thus, the encoding determines how transferable and reusable the learned representations are across different but related tasks. In many approaches, process instances are transformed into fixed-size feature vectors labeled with a prediction target, as required in supervised ML settings such as classification or regression (Di Francescomarino and Ghidini 2022). However, especially in DL-based models, alternative representations such as event sequences or graph-structured inputs are also employed, allowing models to process richer structural information directly (Weinzierl 2022). Typically, two essential encoding steps are necessary: first, sequence encoding, to capture the sequential structure of the traces, and second, event encoding, to encode the information contained in the events (i.e., the event attributes). For the sequence encoding, various strategies have been proposed. These strategies range from simple boolean encodings indicating the presence of specific activities to more detailed representations that preserve the full control flow by including the complete sequence of events or the last $n$ events (states) (Di Francescomarino and Ghidini 2022). The encoding of the event attributes depends largely on their data types. Categorical attributes, such as activity names or resources, are either one-hot encoded or represented by an embedding vector, that is, a dense vector capturing the semantics. Numerical attributes usually require only normalization or scaling into a particular interval (Teinemaa et al. 2019). Temporal information, especially timestamp attributes, can be encoded by deriving features, such as time since the last event or the weekday, which then can be treated as numerical or categorical values (Tax et al. 2017). Additional data attributes on the process context can be present in the form of text. For these attributes, simple natural language processing techniques can be used, such as bag-of-n-grams or latent Dirichlet allocation (Teinemaa et al. 2016), to extract semantically shallow features from text data. By using more advanced techniques that are based on (large) language models, such as contextualized word embedding models like bidirectional encoder representations from transformers (BERT) or robustly optimized BERT pretraining approach (RoBERTa), semantically rich features can be extracted from text data Cabrera et al. (2022); Liessmann et al. (2024).
@@ 2.3 Transfer learning
Typical ML-based approaches, like the ML-based PPM techniques described in Section 2.2, require the availability of labeled data with the same distribution in training and test sets (Pan and Yang 2009). This might not be the case in practice. TL offers a solution, as it “allows the domains, tasks, and distributions used in training and testing to be different” (Pan and Yang 2009, p. 1346). A domain $\pazocal{D}=\{\pazocal{X},P(X)\}$ is described by the feature space $\pazocal{X}$, that is, the set of all term vectors, and the marginal probability distribution $P(X)$, where $X$ is an instance set $X=\left\{\mathbf{x}\mid\mathbf{x}_{i}\in\pazocal{X},\text{with }i=1,\dots,n\right\}$. Task $\pazocal{T}$ consists of the label space $\pazocal{Y}$, that is, the set of all labels, and the predictive function $f(\cdot)$. In TL, a prediction model for the domain with limited data, that is, the target domain, can be improved by transferring knowledge from another domain, the source domain (Weiss et al. 2016). More specifically, the learning of the target prediction function $f_{T}(\cdot)$ for the target learning task $\pazocal{T}_{T}$ and target domain $\pazocal{D}_{T}$ can be improved by transferring knowledge from the source domain $\pazocal{D}_{S}$ and source task $\pazocal{T}_{S}$ with either $\pazocal{D}_{S}\neq\pazocal{D}_{T}$ or $\pazocal{T}_{S}\neq\pazocal{T}_{T}$ (Pan and Yang 2009).
In the context of PPM, a domain is represented by the encoded event data $X$. The feature space $\pazocal{X}$ of the encoded event data $X$ is determined by the selected event attributes, the type of sequence encoding, and the type of event encoding. The marginal distribution $P(X)$ of the encoded event data $X$ is the distribution of the features used to represent the event data $X$. The addressed PPM task (e.g., the prediction of a process-related outcome) is denoted by $\pazocal{T}$. Its goal is to learn a prediction function $f(\cdot)$ that maps a running process instance to a label $y\in\pazocal{Y}$, where $y$ is a value and $\pazocal{Y}$ is the space of all possible values, respectively, for the PPM task $\pazocal{T}$.
@@
The field can be categorized from a problem and a solution perspective, as shown in Figure 2 (Zhuang et al. 2020). From the problem perspective, depending on the existence of the label, three categories exist: i) transductive TL (label only present in $\pazocal{D}_{S}$, $\pazocal{D}_{S}\neq\pazocal{D}_{T}$, $\pazocal{T}_{S}=\pazocal{T}_{T}$), ii) inductive TL (label present in $\pazocal{D}_{T}$, $\pazocal{T}_{S}\neq\pazocal{T}_{T}$), and iii) unsupervised TL (label unknown in both domains, $\pazocal{T}_{S}\neq\pazocal{T}_{T}$) (Pan and Yang 2009). Another differentiation can be made by comparing the feature spaces of the source and target domains. While homogeneous TL refers to $\pazocal{X}^{S}=\pazocal{X}^{T}$, in heterogeneous TL, the feature spaces of the source and the target domains are not the same ($\pazocal{X}^{S}\neq\pazocal{X}^{T}$) (Weiss et al. 2016). From a solution perspective, Pan and Yang (2009) suggest four categories of approaches based on what is transferred: i) instance-based, ii) feature-based, iii) parameter-based, and iv) relational-based. In instance-based approaches, instances from the source domain are re-weighted to be used in the target domain. Feature-based approaches aim to transfer knowledge with a learned feature representation. This can be done asymmetrically, meaning transforming source features to align with target features, or symmetrically, which means discovering a common new feature space (Zhuang et al. 2020). In parameter-based approaches, shared (hyper-)parameters or hyperpriors between the source and target domains are transferred. Lastly, relational-based approaches transfer knowledge in the form of relations between the source and target data, assuming that the data are not independent and identically distributed (Pan and Yang 2009).

### v2

@@
Event logs reflect the behavior of business processes that are mapped in organizational information systems. Predictive process monitoring (PPM) transforms these data into value by creating process-related predictions that provide the insights required for proactive interventions at process runtime. Existing PPM techniques require sufficient amounts of event data or other relevant resources that might not be readily available, which prevents some organizations from utilizing PPM. The transfer learning-based PPM technique presented in this paper allows organizations without suitable event data or other relevant resources to implement PPM for effective decision support. This technique is instantiated in both a real-life intra- and an inter-organizational use case, based on which numerical experiments are performed using event logs for IT service management processes. The results of the experiments suggest that knowledge of one business process can be transferred to a similar business process in the same or a different organization to enable effective PPM in the target context. The proposed technique allows organizations to benefit from transfer learning in intra- and inter-organizational settings by transferring resources such as pre-trained models within and across organizational boundaries.
@@ 1 Introduction
Business processes are the backbone of organizational value creation (Dumas et al. 2018). With ongoing digitalization, business processes are increasingly mapped in information systems (IS), such as enterprise resource planning (ERP) systems, and executed by process users and other stakeholders (Beverungen et al. 2021). Consequently, business processes leave behind large amounts of digital event data that reflect their behavior (van der Aalst 2016). To transform these data into value, various business process analytics approaches have been developed (zur Muehlen and Shapiro 2009).
One of these approaches is predictive process monitoring (PPM) – a set of techniques that leverage historical event log data (Maggi et al. 2014) to create prediction models (Grigori et al. 2004), capable of predicting various targets in running business process instances, such as the next activity, remaining time, or process-related outcomes (Di Francescomarino et al. 2018). By creating these process-related predictions, PPM enables proactive interventions in running business process instances, helping to reduce risks and optimize resource allocation (Di Francescomarino et al. 2018). This leads to improved business process performance (Marquez-Chamorro et al. 2017), even across organizational boundaries (Oberdorf et al. 2023). Therefore, PPM is gaining momentum in business process management (BPM), while its techniques are considered a new class of process mining techniques (Di Francescomarino and Ghidini 2022).
Most PPM techniques rely on shallow machine learning (ML) models (Marquez-Chamorro et al. 2017) or deep learning (DL) models (Rama-Maneiro et al. 2023) to generalize knowledge from historical to unseen event log data (Weinzierl et al. 2024). These models are trained using historical event log data and then applied to new instances of a running business process in order to predict various targets depending on the business context and the associated analytical goals (Maggi et al. 2014). However, organizations – or different departments or subsidiaries within the same organization – often lack the prerequisites for building effective prediction models. These include access to sufficient amounts of event log data (Käppel et al. 2021) and the technical infrastructure required to train complex models, such as graphical processing units (arpteg2018software; Rama-Maneiro et al. 2023).
@@
The field of transfer learning (TL) provides techniques specifically designed to support such scenarios (Pan and Yang 2009), enabling the transfer of ML models and the knowledge they contain from one learning problem (i.e., domain and task) to another. In PPM, TL can be applied in intra-organizational scenarios, in which a prediction model is transferred from one process to another within a single organization (e.g., across departments or subsidiaries), and in inter-organizational scenarios, where prediction models are transferred between processes of different organizations. In both scenarios, we consider all resources to be transferred as process knowledge. This includes the prediction model (also called the base model). Process knowledge originates in its source context, which includes the business process itself, the event data produced by it, and the specific PPM task being addressed using the event data of the business process. The transferred process knowledge is then applied in the target context (see Figure 1).
Figure 1: Transfer process knowledge from source to target context in PPM.
Building on this idea, we present a TL-based technique for PPM that relies on existing similarities between business processes and aims to apply a prediction model to a target context. Therefore, the prediction model is transferred together with further relevant resources, such as information on the corresponding encoding strategy. In particular, this work contributes to the existing body of knowledge in four ways:
@@
    We present a TL-based technique for PPM that enables the transfer of a base model for an outcome-oriented prediction target without the necessity to fine-tune it on the event data from the target context. Our approach also provides guidance on which resources need to be transferred, including pre-trained models, information on the corresponding encoding strategy, further preprocessing details, and metrics used for measuring the prediction performance.
@@
    We propose a cross-domain encoding strategy for event data that enables effective knowledge transfer from a source to a target context. While timestamp information is encoded via a relative mapping approach, activity information is encoded using pre-trained embedding models to preserve semantic consistency across contexts.
@@
    We empirically evaluate our TL-based PPM technique in an intra- as well as an inter-organizational ITSM use case. The results show that the transferred base model achieves a clearly higher prediction performance than models built with traditional PPM techniques on the event data of the target context.
To carry out our research, we adopt the computational design science paradigm (Rai et al. 2017) and structure our work according to the design science research (DSR) publication schema by Gregor and Hevner (2013): First, we present relevant background on ML and DL in PPM, and TL in general (Section 2). Second, we explain the practical relevance of TL for PPM in organizations and introduce inter- and intra-organizational TL in the context of PPM (Section 3). Subsequently, we give an overview of related work on TL in PPM and position our proposed technique with regard to it (Section 4). We then provide a detailed description of the proposed artifact (Section 5). Following this, the evaluation of our artifact is described by introducing the intra- and inter-organizational use case and providing details on the setup used in the evaluation (Section 6). Finally, we present the results of our evaluation (Section 7), discuss implications and future research directions (Section 8), and conclude the paper (Section 9).
@@ 2.1 Event log data and predictive process monitoring
A business process is defined as a sequence of interrelated activities and decisions undertaken to achieve a valuable outcome for a customer (Dumas et al. 2018). Each execution of a business process is referred to as a process instance or case (van der Aalst 2016). Information technology (IT) systems (e.g., ERP systems) commonly record these executions in the form of event logs, that is, collections of timestamped events described by various event attributes that store information about the execution of activities (i.e., steps in the process) (van der Aalst 2016, p. 130). Every event contains at least a case identifier, the name of the executed activity, and a timestamp indicating when the event occurred. Events belonging to the same process instance can be temporally ordered by their timestamp to form a so-called trace (van der Aalst 2016, p. 134).
@@ 2.2 Machine learning in predictive process monitoring
A branch of early PPM approaches (e.g., van der Aalst et al. 2011; Rogge-Solti and Weske 2015) create prediction models by augmenting given or discovered process models with predictive capabilities (Oberdorf et al. 2023). These PPM approaches are called process-aware because they use prediction models that exploit an explicit representation of a process model (e.g., a probabilistic finite automaton or a stochastic Petri net) to make a prediction (Marquez-Chamorro et al. 2017). However, as real-world processes are usually more complex than reflected in process models (van der Aalst 2011), the process-model dependence limits the generalizability of the prediction models. To overcome this limitation, a more recent branch of PPM approaches (e.g., Maggi et al. 2014; Tax et al. 2017) proposes encoding an event log’s traces as feature vectors for the straightforward building of models using ML algorithms (Oberdorf et al. 2023). These approaches are called non-process-aware as the ML models representing the prediction models do not use an explicit description of the process model (Marquez-Chamorro et al. 2017). By leveraging the generalization capabilities of ML models, the approaches often achieve a higher prediction performance than prediction models built on top of process models (Marquez-Chamorro et al. 2017).
@@
Besides a wide range of conventional ML algorithms such as linear regressions, logistic regressions, decision trees, random forests, and gradient boosted decision trees, DL algorithms in the form of deep neural networks (DNNs) have been applied for building prediction models (Marquez-Chamorro et al. 2017; Teinemaa et al. 2019; Verenich et al. 2019). They have gained prominence due to their ability to learn complex patterns and dependencies in large datasets (LeCun et al. 2015). Also in PPM, techniques relying on DNNs have shown superior prediction performance and generalizability across various PPM tasks (Kratsch et al. 2021). Therefore, DNNs have become the state-of-the-art in PPM for building prediction models.
@@
In recent years, a variety of neural network architectures have been employed for building models for different prediction targets, including multi-layer perceptrons (MLPs), convolutional neural networks (CNNs), and recurrent neural networks (RNNs), with either LSTM cells or gated recurrent units (GRUs) (Neu et al. 2022; Rama-Maneiro et al. 2023; Weinzierl et al. 2024). Among these, RNNs, particularly LSTM networks, have been the most widely adopted in PPM. For example, Evermann et al. (2017) apply LSTM models for the next activity prediction, whereas Tax et al. (2017) use LSTM models to predict both the next activity and its corresponding timestamp. More recently, transformer-based models, originally introduced in natural language processing, have shown high prediction performance. These models utilize self-attention mechanisms to effectively capture long-range dependencies in sequential data, as present in event data. For instance, Zaharah A. Bukhsh (2021) proposed a transformer model capable of predicting next activities, remaining time, and timestamps. Beyond efforts to enhance prediction performance through increasingly sophisticated and powerful neural network architectures, complementary strategies have focused on enriching the input data by incorporating additional sources of information. While many prediction models primarily rely on control-flow information and time information derived from timestamps (e.g., elapsed time), several approaches extend this by integrating contextual information, for example, in the form of case-level or resource attributes (Weinzierl et al. 2019). For example, Di Francescomarino et al. (2018) propose multi-perspective models that combine control-flow and contextual information for outcome prediction. As another example, Hinkka et al. (2019) suggest clustering categorical event attributes and using the resulting cluster labels as additional information for a RNN. In contrast to these approaches, Pegoraro et al. (2021) and Cabrera et al. (2022) integrate unstructured data in the form of text as contextual information. The ability of DL methods to handle diverse prediction tasks and to incorporate heterogeneous data sources not only demonstrates their flexibility but also hints at the potential for reusing learned representations in varying contexts (Weinzierl et al. 2024).
@@
To build a prediction model using an ML or DL algorithm, it is crucial to appropriately encode the underlying event data, as the quality and structure of the encoding strongly affect the model’s ability to generalize. Thus, the encoding determines how transferable and reusable the learned representations are across different but related tasks. In many approaches, process instances are transformed into fixed-size feature vectors labeled with a prediction target, as required in supervised ML settings, such as classification or regression (Di Francescomarino and Ghidini 2022). However, especially in DL-based models, alternative representations such as event sequences or graph-structured inputs are also employed, allowing models to process richer structural information directly (Weinzierl 2022). Typically, two essential encoding steps are necessary: first, sequence encoding, to capture the sequential structure of the traces, and second, event encoding, to encode the information contained in the events (i.e., the event attributes). For the sequence encoding, various strategies have been proposed. These strategies range from simple boolean encodings, indicating the presence of specific activities, to more detailed representations, preserving the full control flow by including the complete sequence of events or the last $n$ events (states) (Di Francescomarino and Ghidini 2022). The encoding of the event attributes depends largely on their data types. Categorical attributes, such as activity names or resources, are either one-hot encoded or represented by an embedding vector, that is, a dense vector capturing the semantics. Numerical attributes usually require only normalization or scaling into a particular interval (Teinemaa et al. 2019). Temporal information, especially timestamp attributes, can be encoded by deriving features, such as time since the last event or the weekday, which then can be treated as numerical or categorical features (Tax et al. 2017). Additional data attributes on the process context can be present in the form of text. For these attributes, simple natural language processing techniques can be used, such as bag-of-n-grams or latent Dirichlet allocation (Teinemaa et al. 2016), to extract semantically shallow features from text data. By using more advanced techniques that are based on (large) language models, such as contextualized word embedding models like bidirectional encoder representations from transformers (BERT) or robustly optimized BERT pretraining approach (RoBERTa), semantically rich features can be extracted from text data (Cabrera et al. 2022; Liessmann et al. 2024).
@@ 2.3 Transfer learning
Typical ML-based approaches, like the ML-based PPM techniques described in Section 2.2, require the availability of labeled data with the same distribution in training and test sets (Pan and Yang 2009). This might not be the case in practice. TL offers a solution, as it “allows the domains, tasks, and distributions used in training and testing to be different” (Pan and Yang 2009, p. 1346). A domain $\pazocal{D}=\{\pazocal{X},P(X)\}$ is described by the feature space $\pazocal{X}$, that is, the set of all term vectors, and the marginal probability distribution $P(X)$, where $X$ is an instance set $\left\{\mathbf{x}\mid\mathbf{x}_{i}\in\pazocal{X},\text{with }i=1,\dots,n\right\}$. Task $\pazocal{T}$ consists of the label space $\pazocal{Y}$, that is, the set of all label values, and the predictive function $f(\cdot)$. In TL, a prediction model for the domain with limited data, that is, the target domain, can be improved by transferring knowledge from another domain, the source domain (Weiss et al. 2016). More specifically, the learning of the target prediction function $f_{T}(\cdot)$, for the target learning task $\pazocal{T}_{T}$, and target domain $\pazocal{D}_{T}$, can be improved by transferring knowledge from the source domain $\pazocal{D}_{S}$, and source task $\pazocal{T}_{S}$, with either $\pazocal{D}_{S}\neq\pazocal{D}_{T}$ or $\pazocal{T}_{S}\neq\pazocal{T}_{T}$ (Pan and Yang 2009).
In the context of PPM, a domain is represented by the encoded event data $X$. The feature space $\pazocal{X}$ of the encoded event data $X$ is determined by the selected event attributes and the used sequence and event coding. The marginal probability distribution $P(X)$ of the encoded event data $X$ is the distribution of the features used to represent the event data $X$. The addressed PPM task (e.g., the prediction of a process-related outcome) is denoted by $\pazocal{T}$. The prediction function $f(\cdot)$ of the task aims to map a running process instance to a label value $y\in\pazocal{Y}$, where $\pazocal{Y}$ is the space of all possible label values for the PPM task $\pazocal{T}$.
@@
The field can be categorized from a problem and a solution perspective, as shown in Figure 2 (Zhuang et al. 2020). From the problem perspective, depending on the existence of the label, three categories exist: i) transductive TL (label only present in $\pazocal{D}_{S}$, $\pazocal{D}_{S}\neq\pazocal{D}_{T}$, $\pazocal{T}_{S}=\pazocal{T}_{T}$), ii) inductive TL (label present in $\pazocal{D}_{T}$, $\pazocal{T}_{S}\neq\pazocal{T}_{T}$), and iii) unsupervised TL (label unknown in both domains, $\pazocal{T}_{S}\neq\pazocal{T}_{T}$) (Pan and Yang 2009). Another differentiation can be made by comparing the feature space of the source and the target domain. While homogeneous TL refers to $\pazocal{X}^{S}=\pazocal{X}^{T}$, in heterogeneous TL, the feature space of the source and the target domain is not the same ($\pazocal{X}^{S}\neq\pazocal{X}^{T}$) (Weiss et al. 2016). From a solution perspective, Pan and Yang (2009) suggest four categories of approaches based on what is transferred: i) instance-based, ii) feature-based, iii) parameter-based, and iv) relational-based. In instance-based approaches, instances from the source domain are re-weighted to be used in the target domain. Feature-based approaches aim to transfer knowledge with a learned feature representation. This can be done asymmetrically, meaning transforming source features to align with target features, or symmetrically, which means discovering a common new feature space (Zhuang et al. 2020). In parameter-based approaches, shared (hyper-)parameters or hyperpriors between the source and the target domain are transferred. Lastly, relational-based approaches transfer knowledge in the form of relations between the source and target data, assuming that the data are not independent and identically distributed (Pan and Yang 2009).